# HPC Alongside User-space Kubernetes

### v1

@@ Abstract
High performance computing (HPC) and cloud have traditionally been separate, and presented in an adversarial light. The conflict arises from disparate beginnings that led to two drastically different cultures, incentive structures, and communities that are now in direct competition with one another for resources, talent, and speed of innovation. With the emergence of converged computing, a new paradigm of computing has entered the space that advocates for bringing together the best of both worlds from a technological and cultural standpoint. This movement has emerged due to economic and practical needs. Emerging heterogeneous, complex scientific workloads that require an orchestration of services, simulation, and reaction to state can no longer be served by traditional HPC paradigms. However, while cloud offers automation, portability, and orchestration, as it stands now it cannot deliver the network performance, fine-grained resource mapping, or scalability that these same simulations require. These novel requirements call for change not just in workflow software or design, but also in the underlying infrastructure to support them. This is one of the goals of converged computing. While the future of traditional HPC and commercial cloud cannot be entirely known, a reasonable approach to take is one that focuses on new models of convergence, and a collaborative mindset. In this paper, we introduce a new paradigm for compute – a traditional HPC workload manager, Flux Framework, running seamlessly with a user-space Kubernetes “Usernetes” to bring a service-oriented, modular, and portable architecture directly to on-premises HPC clusters. We present experiments that assess HPC application performance and networking between the environments, and provide a reproducible setup for the larger community to do exactly that.
@@ I Introduction
Modern paradigms for computing are in a phase of transition. While high performance computing (HPC) has traditionally been the powerhouse for scientific workflows, Kubernetes is taking over the space as a new economic leader, having a projected annual growth rate (CAGR) of 20% by 2027 to compare to HPC’s much smaller rate of 8% by 2026. This is a valuation of over 1.1 trillion dollars for cloud as compared to 40 billion for HPC [1, 2, 3]. The influx of money into the cloud easily offers promise for innovation to further drive demand, allowing for features that include modularity, portability, and automation of workflows. HPC is presented with the same problem as any organism in a changing environment: it needs to adapt or face a potential decline. The numbers tell the story [4] with HPC being oriented to non recurring engineering costs (NREs) that mean a loss of money with each new system, and cloud being positioned to lead the innovation space with an influx of money. However, while Kubernetes and associated cloud technologies are clear leaders in the speed of development and adoption, there is much to be learned from the rich history and technical expertise of the HPC community. It would be a mistake to not push for approaches that enable these two communities to work together, learn from one another, and unify traditionally competitive or disparate technologies.
Paradigms of computing have evolved over time, and in a way that is supportive of the needs of its users. Whether supporting scientific domains, medicine, advanced engineering, or business, the fundamental ingredients of a system typically come down to a choice of physical hardware for compute and storage, operating systems, software and methods to provision it, and administration of resources and user permissions. HPC, by way of supercomputers, is such a design, experiencing an unbelievable range of growth in a span of decades to go from processing 1k operations per second to, in present day, over 100 quadrillion [5]. As the needs of scientific discovery were at the forefront of incentives, supercomputers have developed at national laboratories, and academic and private institutions as powerhouses of compute. The need to support multiple uses on these massive systems has led to models of multi-tenancy, and the need for maximally utilizing the resources has led to the emergence of programming models, architectures, and hardware, often selected strategically for custom systems [6, 7, 8]. While a skeptic might look at HPC technologies as compared to the innovation of cloud and consider them now outdated, if it weren’t for HPC pushing for the early development of hardware, communication protocol, operating systems, networking, and parallel applications, commercial technologies would not be where they are today. While HPC has continued to optimize for niche scientific use cases that require massive scale and performance, commercial cloud has taken these same technologies and optimized for generalizability, an opposing incentive structure that has led to drastically different economic outcomes.
The incentive structure of HPC to be optimized for specific scientific use cases has led to many niche, “snowflake” setups that make applications less portable and systems more unique. The culture of the HPC community has followed suit. With a focus on scale and performance, annual competitions such as the Top500 [9] encourage building new systems that are more powerful and impressive than previous attempts. With an economic market eager to invest in the newest, and fastest hardware, vendors followed, including AMD, NVIDIA, Intel, IBM, and HPE, among others [10], and often creating markets where a small number of vendors retained the majority of market share [11]. The resulting landscape prioritized demonstrations of power and performance, and the HPC community was less impacted by, for example, the reproducibility crisis [12, 13, 14] in the mid 2010s that confronted specific scientific disciplines and challenged their ability to do sound, reproducible work [15]. Arguably, this inflection point created a slight separation in incentives between some of the scientific community that prioritized reproducible, portable environments for running pipelines (e.g., the biosciences) and the HPC community that served them. Centers that needed to run physics or climate simulations at scale arguably still cared about performance, often with reproducibility or portability as an afterthought. When commercial cloud emerged as an option for compute, even for research groups, the modularity that it offered was attractive to scientists and small and medium businesses alike. Especially for this subset of scientific communities that ran embarrassingly parallel pipelines that often didn’t require HPC programming models or performant networks the cloud was highly attractive, and workflow tools were developed that harnessed cloud virtual machines and eventually container technologies [16, 17] as early as 2011. These pipelines had a focus on the portability aspect. Container technologies did not come to HPC until much later [18, 19, 20], and well after the peak of the reproducibility crisis as researchers began to discover and request them for use on their systems. From the above, it can be suggested that the desire and underlying incentive structure to want cloud native features, namely portability, modularity, and automation, came to the HPC community with some delay, and this is reflected in the underlying systems that are not designed to support them. The reasons likely come back to incentives. Whereas commercial cloud optimized for general portability, HPC was still driven by performance, and the real need to push the scale of simulations. However, as machines started to become more heterogeneous, a gap emerged between the system theoretical deliverable FLOPs, and the FLOPs usable by many simulation-driven HPC applications.
Cloud computing, the arm of compute that emerged onto the scene in 2006 with Amazon’s first commercial cloud [21], embodied many of these features by way of offering compute as services. The design also reflects the needs of the customers – isolation of resources such as storage, compute, and networking was essential for privacy, cost containment, and ownership of data. It’s interesting to consider how early cloud computing was presented from a branding standpoint, because early reports often portrayed cloud as a series of services, and often to supplement on premises systems [22, 23]. This meant that early on, cloud was not considered a competitor to HPC because it lacked use cases beyond simple services, and did not perform well. HPC centers likely did not view cloud as a serious competitor in the space of hiring or control of vendor production. Based on the current landscape, where cloud is the primary driver of economic innovation and regularly steals talent from the HPC community, at some point this had to change. It could be argued that this change started with the “big data” trend in the early 2010s, where companies could make money off of data analysis, allowing the eventual progression into support for AI/ML. It might also be suggested that HPC has typically been a resource of privileged access, typically only available to researchers at a specific institution, and at best more publicly available but still only with the proper identity or registration step. This left open a hole in the market – computing resources for profit-driven startups and companies that could not afford their own systems and did not quality for academic, HPC ones – that cloud vendors eventually filled. While HPC aimed to maximize science and fine tune resources for specific, niche scientific problems of the few, commercial cloud aimed to maximize profits and generalize resources for computational problems of the many. However, a new competitor in the space was not enough to push for entirely new paradigms of compute. Arguably, these two communities remained disparate for several years, because the user bases, scientists vs. small groups without access to HPC, remained fairly separate. However, this separation is going away because of a new, common underlying unit: the applications themselves.
@@
Applications and workflows are changing, becoming more complex and heterogeneous to reflect the expanding set of techniques that are used to conduct science. The original models common for HPC, specifically models that were embarrassingly parallel [24] or distributed [25, 26] postulated many tasks running at scale. More modern scientific workflows extend beyond this simple design, often requiring multiple steps that are represented in a directed acyclic graph (DAG) and sent to a workload manager with a queue and scheduler to be provisioned resources. This task falls under the jurisdiction of a workflow tool, and increasingly the units of work required to complete the entire workload include services such as databases and task queues, steps with multiple components working together, and temporal variability that can vary from almost instantaneous execution to training that might last days, weeks, or longer.
@@
In the last few years machine learning (ML) and artificial intelligence (AI) has come to the forefront of interest for industry and science. While the bread and butter of HPC applications has traditionally been parallel programming models such as the message passing interface (MPI), a separate stream of tools for ML have emerged that don’t always use MPI, but instead reflect the architecture on which they need to run. For example, ML training provided by software like Pytorch or Tensorflow that requires matrix operations has developed to take advantage of single large machines running GPUs [28]. Interestingly, the development has come full circle – with these same models and matrix operations being architected to again fit HPC systems, but extended to use both GPU and MPI to scale across physical machines. Another design pattern that has emerged for ML, and specifically applications that require a distributed architecture, is a set of machines that adhere to a leader-worker design [29] where a single machine allocates units of work that can be expanded or contracted, restarted on failure, and generally fault tolerant. This is in contrast to MPI, which is not fully fault tolerant nor elastic, except for approaches that attempt to recover failures [30] or restart an entire component of the application [31]. The trade-off between the tightly integrated but more fragile MPI and distributed, fault tolerant and loosely integrated worker design is performance. A decision to use one technology or the other, or even choose on-premises HPC vs. cloud environment, often comes down to a choice on where to operate along this spectrum. The users of high performance computing systems have the desire for these features, and it is often the case that the infrastructure and tooling simply cannot support them.
@@ II-A1 Kubernetes
Kubernetes is a resource management and scheduling powerhouse of industry, with 71% usage reported by Fortune 500 companies and over 74K contributors [40, 41]. It has become the de-facto standard tool for container orchestration, offering a developer and application environment that is modular, configurable, and can be deployed on demand. Historically, the barrier for adoption of such an infrastructure on shared systems was the need for root daemon services, however with rootless technologies, that has changed, offering the first opportunity for the HPC community to explore the technology space.
@@ II-A2 Usernetes
Usernetes grew out of the concept of an unprivileged or “rootless” container, which was first introduced by LXC in 2014 [42]. User namespaces were available since version 3.12 of the Linux kernel, and they provide a large range of user and group identifiers that can be intelligently mapped into the container, both for access permissions and resources such as network, processes, and devices. This meant that, unlike a root daemon-run container such as the original version of Docker, escaping from the container would not give the attacker root access to a system. The container runtime “runc” followed suit, adding support for rootless in 2017 [43] and this was quickly followed by major container technologies such as Docker and Podman, among others [44]. It very naturally followed that rootless containers could be used in Kubernetes by way of first adding features to the kubelet [45, 46] and then developing the full Usernetes project [47] in 2018. Generation 1 of the project, which persisted through early 2023, was a collection of bash scripts that did not require a rootless container technology, and did not have a multi-host use case in mind. One of the authors of this paper (VS) pursued this multi-host use case in early 2023, and started engagement with the project [48]. Generation 2 was released in the same year [49] that harnessed fully rootless container technologies to deploy user-space Kubernetes, and support this use case. This development provided a new opportunity to deploy an entire Kubernetes in user-space, and alongside a traditional HPC workload manager, which was the vision of the authors of this paper at the time. Getting to the point of having a fully working environment, however, would prove to be challenging.
@@ II-B1 Mitigating Resource Management Conflict
The Flux Framework provides an alternative approach: hierarchical scheduling and resource management [39]. Flux is based on the principles of hierarchical bounding, meaning that a parent grants resources to its children, and instance effectiveness, signifying that each instance can be independent and is solely responsible for its resource set [51]. A hierarchy of schedulers and resource managers does not suffer from the same throughput limitations as a monolithic scheduler: a hierarchy of Flux instances can schedule over 100 million jobs in a single workflow [52] and achieve nearly 800 jobs per second scheduling throughput [53]. Deploying a second resource manager and scheduler within a Flux hierarchy with instance effectiveness bounds the scope of scheduling and resource management conflict to the instance level of deployment. Thus a second resource manager and scheduler like Usernetes can run in an instance encapsulated as a job in a hierarchy, providing lifetime control, fairshare, user-scoped utilization data, and fine-grained resource management mapping that are absent or poorly supported in Usernetes and Kubernetes. Currently, Flux only supports static job resource allocations, hence a level of a Flux hierarchy consisting of Usernetes and another Flux instance is a static partitioning of resource managers. However, the maximum resource fragmentation is bounded by the number of resources managed by the containing Flux instance in the hierarchy. Adding support for allocation dynamism can reduce the upper bound of fragmentation. At that point mitigating conflict between dynamic Usernetes and Flux allocations will be necessary. Adding Flux Framework support for allocation elasticity and resource conflicts arising from the elasticity is planned future work.
@@ II-B2 System Requirements
Generation 2 of Usernetes uses the base image provided by Kubernetes in Docker “Kind” [54] project, and uses kubeadm for cluster configuration [55], containerd for the container runtime interface (CRI) [56], runc for the OCI-standardized container spawning [57], and Flannel as the network fabric for the container networking interface (CNI) [58]. Supported operating systems at the time of this work included Ubuntu 22.04 (recommended), along with Rocky Linux and Alma Linux 9. A rootless container technology is required, with Docker as first preference, but Podman and nerdctl also being supported. To support rootless mode, cgroup version 2 delegation must be enabled [59], allowing for consumption of cpu, memory, I/O and the process namespace. The br_netfilter [60] and vxlan [61] network modules must be enabled to allow for inter-node communication, and network packet checking set to a less stringent mode [62]. One known weakness in the setup is the use of slirp4netns for network traffic, which receives packets from a container via a TAP device and extracts payloads to send via a socket [63]. In practice, this intermediate relay of packets leads to a slowdown in network transmission, and the authors of this paper had seen workloads to be at least twice as slow using traditional ethernet based on early experimentation [64]. These system-level, often security relevant changes introduce a barrier to quick prototyping on an existing bare-metal system, as it either is the case that changes need to be thoroughly vetted, or it simply is not possible for a developer with user-level permissions. A more reasonable development approach would be to deploy to a testing environment with virtual machines.
@@ II-B3 Deployment
The choice of where to deploy one or more virtual machines that are oriented to run Usernetes also presents considerable challenges, as each potential environment can have differences with respect to nuances of networking, kernel features available, or namespaces. The authors of this paper first attempted several cloud setups, and failed due to issues with controlling or exposure of network resources. The next avenue to pursue, which was a success, was prototyping a setup of an HPC workload manager running alongside Usernetes in Linux Virtual Machines “LIMA” [65]. This was followed by a small, private on-premises cluster deployed with Ovirt [66] that provided the first performance testing that was presented at FOSDEM in 2024 [64].
@@
Amazon Web Services (AWS) Elastic Compute (EC2) was used to deploy a set of 33 virtual machines, a family of instances “hpc7g” [67] that are powered by AWS Graviton (ARM) processes, oriented toward HPC workloads in that they offer, by default, single threading, exclusive access to the full virtual machine, and a reasonable network (128 GiB memory, and 200 Gbps EFA Network bandwidth). While the machine is offered in three sizes (16, 32, and 64, physical cores, respectively) because the memory does not change, this means choosing between a smaller number of cores that each have more memory, or more cores with less memory. For the purposes of this work, the authors chose the smallest size, hpc7g.4xlarge, choosing for the highest ratio of memory per core for the performance benefits [68]. A cluster placement group was used to ensure selection of nodes that are physically close together [69]. The Elastic Fiber Adapter (EFA) network interface [70] was essential for the basic functioning of MPI-dependent HPC applications in this setup on EC2. Such an interface that bypasses the operating system and goes directly to a driver provides the minimal application performance needed for communication protocols such as MPI, and bypasses some of the problems of using slirp4netns. Without such an interface, many of these workloads simply don’t run, or run so unreasonably slow it would be impossible to achieve any kind of result in a reasonable amount of time.
There were several key features or workarounds specific to AWS that ultimately afforded the final setup to work. Firstly, AWS does not provide an easy means to generate predictable hostnames for instances unless a Route 53 owned hostname is available [71]. This means that a configuration file for Flux Framework, in which all the node hostnames are known, is not possible to generate apriori. The workaround for this problem is to have a startup script that uses the AWS Python client [72] to query for the set of instances via a selector label, and then compare the number found to a count (derived programmatically from the deployment configuration) to wait for instances to be online, retrieve their hostnames, and dynamically write the broker configuration file. This startup script is also useful for impromptu changes that need to be applied to all virtual machines such as system configuration or additional software installations. Secondly, default health check periods presented difficulties during testing. In particular, an instance would be marked or found to be unhealthy, taken offline to be replaced, and the running experiment would fail. Depending on the instance taken offline, an application run may fail or the virtual cluster itself can fail, requiring manual reset. Setting this health check value to a high value (25 hours for our experiments) resolved this issue. Thirdly, the network setup was challenging. In particular, AWS services such as the load balancer that provides the ip addresses for the bootstrapping of the workload manager network typically require subnets in two or more availability zones. This presented a problem to Usernetes, as internal ip addresses that it required could not be seen across zones, and the workers could not register to the control plane. The solution to the problem was to deploy with two or more subnets and availability zones to pass validation, but designate the instance group to use only one. Finally, the Elastic Fiber Adapter requires the security group to have an ingress and egress rule to itself, which is not clearly documented, nor is it an intuitive or obvious thing to do.

### v2

# A Contention-Free Model for Converged Kubernetes on HPC

@@ Abstract
Next-generation heterogeneous, complex scientific workloads that require integration of service orchestration, simulation, and automated management of state are not well supported by traditional high performance computing (HPC). While alternative cloud options offer automation, portability, and orchestration, cloud does not generally deliver the network latency performance, fine-grained resource mapping, or application scalability that tightly-coupled HPC simulations can require. These novel requirements call for change in workflow software or design and in the underlying supporting infrastructure. Converged computing is an emerging paradigm that represents the best of both worlds of cloud and high performance computing (HPC) technology and culture, and such a paradigm offers a solution to this predicament. In this paper, we introduce a new model for compute resource co-management—an HPC workload manager, Flux Framework, running seamlessly with user-space Kubernetes “Usernetes” to instantiate an automated, service-supporting, modular, and portable converged architecture to on-premises HPC clusters. We present the first setup of this kind, and experiments that evaluate HPC application and network performance, finding equivalent performance for bandwidth, and an MPI application that heavily utilizes CPU. We provide a reproducible experimental and deployment setup to enable the larger community to utilize and further collaborate on the work.
@@ I Introduction
The landscape of high performance computing (HPC) is changing. Applications and workflows are becoming more complex and heterogeneous to reflect the expanding set of techniques that are used to conduct science. Early parallel computing in HPC was exemplified by embarrassingly parallel [1], bulk synchronous parallel [2], or distributed [3, 4] models. Reflecting the increasing interdisciplinary nature of computational science, modern scientific workflows extend beyond early designs to integrate pipelined simulations, ensemble methods, coupled multiscale models, in-situ visualization, AI/ML training or surrogate model inference, data movement and staging, etc. The components can be steps represented in a directed acyclic graph (DAG) and sent to a workload manager with a queue and scheduler for resource allocation and task management. Representing the steps in a DAG enables encoding complex task relationships into a well-defined data structure. This task falls under the jurisdiction of a workflow tool, and increasingly the units of work required to complete the entire workload include services such as databases and task queues, steps with multiple components working together, and time scales that can vary from millisecond execution to AI/ML training that can take weeks or more.
@@
In the last few years machine learning (ML) and artificial intelligence (AI) has come to the forefront of interest for industry and science. While the bread and butter of HPC applications has traditionally been parallel programming models such as the message passing interface (MPI), a separate stream of tools for ML have emerged that don’t always use MPI, but instead reflect the architecture on which they need to run. For example, ML training provided by software like Pytorch or Tensorflow that requires matrix operations has developed to take advantage of single large machines running GPUs [6]. Interestingly, the development has come full circle – with these same models being ported back to HPC systems, but extended to use both GPU and MPI to scale across physical machines. Another design pattern that has emerged for ML, and specifically applications that require a distributed architecture, is a set of machines that adhere to a leader-worker design [7] where a single machine allocates units of work that can be expanded or contracted, restarted on failure, and generally fault tolerant. This is in contrast to MPI, which is not fully fault tolerant nor elastic, except for approaches that attempt to recover failures [8] or restart an entire component of the application [9]. The trade-off between the tightly integrated but more fragile MPI and distributed, fault tolerant and loosely integrated worker design is performance. A decision to use one technology or the other, or even choose on-premises HPC vs. cloud environment, often comes down to a choice on where to operate along this spectrum. The users of HPC systems have the desire for all these features, and it is often the case that the infrastructure and tooling simply cannot support them.
@@ II-A1 Kubernetes
Kubernetes is a resource management and scheduling powerhouse of industry, with 71% usage reported by Fortune 500 companies and over 90K contributors [20, 21]. It has become the de-facto standard tool for container orchestration, offering a developer and application environment that is modular, configurable, and can be deployed on demand. Historically, the barrier for adoption of such an infrastructure on shared systems was the need for root daemon services, however with rootless technologies, that has changed, offering the first opportunity for the HPC community to explore the technology space.
@@ II-A2 Usernetes
Usernetes grew out of the concept of an unprivileged or “rootless” container, which was first introduced by LXC in 2014 [22]. User namespaces were available since version 3.12 of the Linux kernel, and they provide a large range of user and group identifiers that can be intelligently mapped into the container, both for access permissions and resources such as network, processes, and devices. This meant that, unlike a root daemon-run container such as the original version of Docker, escaping from the container would not give the attacker root access to a system. The container runtime “runc” followed suit, adding support for rootless in 2017 [23] and this was quickly followed by major container technologies such as Docker and Podman, among others [24]. It followed that rootless containers could be used in Kubernetes by way of first adding features to the kubelet [25, 26] and then developing the full Usernetes project [27] in 2018. Generation 1 of the project, which persisted through early 2023, was a collection of bash scripts that did not require a rootless container technology, and did not support multiple hosts. One of the authors of this paper (VS) pursued this multi-host use case in early 2023, and started engagement with the project [28]. Generation 2 was released in the same year [29] that harnessed fully rootless container technologies to deploy user-space Kubernetes. This development provided a new opportunity to deploy an entire Kubernetes in user-space, and alongside a traditional HPC workload manager.
@@ II-B1 Mitigating Resource Management Conflict
The Flux Framework provides an alternative approach: hierarchical scheduling and resource management [19]. Flux is based on the principles of hierarchical bounding, meaning that a parent grants resources to its children, and instance effectiveness, signifying that each instance can be independent and is solely responsible for its resource set [31]. A hierarchy of schedulers and resource managers does not suffer from the same throughput limitations as a monolithic scheduler: a hierarchy of Flux instances can schedule over 100 million jobs in a single workflow [32] and achieve nearly 800 jobs per second scheduling throughput [33].
Deploying a second resource manager and scheduler within a Flux hierarchy with instance effectiveness bounds the scope of scheduling and resource management conflict to the instance level of deployment. Thus a second resource manager and scheduler like Usernetes can run in an instance encapsulated as a job in a hierarchy, providing lifetime control, user compartmentalization, fairshare, user-scoped utilization data, and fine-grained resource management mapping that are absent or poorly supported in Usernetes and Kubernetes.
@@ II-B2 System Requirements
Generation 2 of Usernetes uses the base image provided by Kubernetes in Docker “Kind” [34] project, and uses kubeadm for cluster configuration [35], containerd for the container runtime interface (CRI) [36], runc for the OCI-standardized container spawning [37], and Flannel as the network fabric for the container networking interface (CNI) [38]. Supported operating systems at the time of this work included Ubuntu 22.04 (recommended), along with Rocky Linux and Alma Linux 9. Usernetes requires a rootless container technology with Docker as first preference, but Podman and nerdctl are also supported. To support rootless mode, cgroup version 2 delegation must be enabled [39], allowing for consumption of CPU, memory, I/O and the process namespace. The br_netfilter [40] and vxlan [41] network modules must be enabled to allow for inter-node communication, and network packet checking set to a less stringent mode [42].
One known weakness in the setup is the use of slirp4netns for network traffic, which receives packets from a container via a TAP device and extracts payloads to send via a socket [43]. In practice, this intermediate relay of packets leads to a slowdown in network transmission, and the authors of this paper observed greater than a 2×2$\times$2 × slowdown using traditional ethernet in early experimentation [44]. These system-level, often security relevant changes introduce a barrier to quick prototyping on an existing bare-metal system, as it either is the case that changes need to be thoroughly vetted, or it simply is not possible for a developer with user-level permissions. A more reasonable development approach was using a testing environment with cloud virtual machines.
@@ II-B3 Deployment
The choice of where to deploy one or more virtual machines that are oriented to run Usernetes also presents considerable challenges, as each potential environment can have differences with respect to nuances of networking, kernel features available, or namespaces. The first successful endeavour was prototyping a setup of an HPC workload manager running alongside Usernetes in Linux Virtual Machines “LIMA” [45]. This was followed by a small, private on-premises cluster deployed with Ovirt [46] that provided the first performance testing that was presented at FOSDEM in 2024 [44].
@@
Amazon Web Services (AWS) Elastic Compute (EC2) was used to deploy a set of 33 virtual machines, a family of instances “hpc7g” (hpc7g.4xlarge) [47] that are powered by AWS Graviton (Arm) processes, oriented toward HPC workloads in that they offer, by default, single threading, exclusive access to the full virtual machine, and a reasonable network (128 GiB memory, and 200 Gbps EFA Network bandwidth). A cluster placement group was used to ensure selection of nodes that are physically close together [48]. The Elastic Fabric Adapter (EFA) network interface [49] was essential for the basic functioning of MPI-dependent HPC applications in this setup on EC2. Such an interface that bypasses the operating system and goes directly to a driver provides the minimal application performance needed for communication protocols such as MPI, and bypasses some of the problems of using slirp4netns. Without such an interface, many of these workloads simply don’t run, or run so unreasonably slow it would be impossible to achieve any kind of result in a reasonable amount of time.
There were several key features or workarounds specific to AWS that ultimately enabled the final setup to work. A startup script using the AWS Python client [50] was used to query the set of instance hostnames via a selector label to populate the Flux configuration file. Secondly, default health checks needed to be disabled via an arbitrary large check duration as an erroneous failed check could bring down the final cluster. Thirdly, the AWS load balancer requiring subnets in two or more availability zones presented a problem to Usernetes that required the addresses but could not interact across zones. The solution was to deploy the setup with two or more subnets and availability zones to pass validation, but designate the instance group to use only one. Finally, the Elastic Fabric Adapter requires the security group to have an ingress and egress rule to itself, which is not clearly documented, nor is it an intuitive or obvious thing to do.